# Quantum key distribution rates from non-symmetric conic optimization

### v1

@@ Abstract
Computing key rates in QKD numerically is essential to unlock more powerful protocols, that use more sophisticated measurement bases or quantum systems of higher dimension. It is a difficult optimization problem, that depends on minimizing a convex non-linear function: the relative entropy. Standard conic optimization techniques have for a long time been unable to handle the relative entropy cone, as it is a non-symmetric cone, and the standard algorithms can only handle symmetric ones. Recently, however, a practical algorithm has been discovered for optimizing over non-symmetric cones, including the relative entropy. Here we adapt this algorithm to the problem of computation of key rates, obtaining an efficient technique for lower bounding them. In comparison to previous techniques it has the advantages of flexibility, ease of use, and above all performance.
@@ 1 Introduction
We use the framework from [winick2018] to formulate the key rate problem as a convex optimization problem, the minimization of a relative entropy over a convex domain. Using the relative entropy cone, which is a non-symmetric cone, it becomes a conic optimization problem. Although for a long time algorithms for optimizing over non-symmetric cones have been known [nesterov1999, tuncel2001, nesterov2012], they were hardly practical, as they required for instance a tractable barrier function for the dual cone. This changed with the discovery of Skajaa and Ye’s algorithm [skajaa2015, papp2017], and with it solving the key rate problem becomes in principle possible. There is, however, a technical difficulty: this formulation of the key rate problem is often not strictly feasible, and such problems cannot be reliably solved with conic optimization methods. The standard technique to deal with this, called facial reduction [drusvyatskiy2017], does apply, but the resulting reduced problem is no longer a minimization of the relative entropy. For this reason we introduce a new convex cone to deal directly with the reduced problem.
@@ 2 QKD rate as a conic problem
$\begin{gathered}\min_{\rho}\ D(\mathcal{G}(\rho)\|\mathcal{Z}(\mathcal{G}(\rho)))\\ \text{s.t.}\quad\rho\succeq 0,\quad\operatorname{tr}(\rho)=1\\ \operatorname{tr}(E_{k}\rho)=p_{k}\ \forall k\end{gathered}$ (3)
@@
$\mathcal{K}_{\text{RE}}=\text{cl}\,\{(h,\rho,\sigma)\in\mathbb{R}\times\mathbb{C}^{n\times n}\times\mathbb{C}^{n\times n};\,\rho\succ 0,\ \sigma\succ 0,\ h\geq D(\rho\|\sigma)\}.$ (4)
It becomes then
$\begin{gathered}\min_{h,\rho}\ h\\ \text{s.t.}\quad\rho\succeq 0,\quad\operatorname{tr}(\rho)=1\\ \operatorname{tr}(E_{k}\rho)=p_{k}\ \forall k\\ \Big{(}h,\mathcal{G}(\rho),\mathcal{Z}(\mathcal{G}(\rho))\Big{)}\in\mathcal{K}_{\text{RE}}\end{gathered}$ (5)
@@
$\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}=\{(h,\sigma)\in\mathbb{R}\times\mathbb{C}^{n\times n};\,\sigma\succeq 0,\ h\geq-H(\widehat{\mathcal{G}}(\sigma))+H(\widehat{\mathcal{Z}}(\sigma))\}.$ (7)
@@
$\begin{gathered}\min_{h,\sigma}\ h\\ \text{s.t.}\quad\operatorname{tr}(\sigma)=1,\quad\operatorname{tr}(F_{k}\sigma)=p_{k}\ \forall k\\ (h,\sigma)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}\end{gathered}$ (8)
@@ 3 Implementation of the QKD cone
In order to successfully optimize over a given convex cone $\mathcal{K}$, both using Skajaa and Ye’s algorithm or traditional methods for symmetric cones, it’s essential to provide a logarithmically homogeneous self-concordant barrier (LHSCB) function for it [nesterov1997]. It is defined as a function $f:\operatorname{int}(\mathcal{K})\to\mathbb{R}$ such that $f(x_{i})\to\infty$ along every sequence converging to the boundary of $\mathcal{K}$, is three times continuously differentiable, strictly convex, and
$f(\tau x)=f(x)-\nu\log\tau\quad\forall x\in\operatorname{int}(\mathcal{K}),\forall\tau>0$ (9)
$|\nabla^{3}f(x)[\xi,\xi,\xi]|\leq 2(\nabla^{2}f(x)[\xi,\xi])^{3/2}\quad\forall x\in\operatorname{int}(\mathcal{K}),\forall\xi$ (10)
@@
$f(h,\rho,\sigma)=-\log(h-D(\rho\|\sigma))-\operatorname{logdet}(\rho)-\operatorname{logdet}(\sigma)$ (11)
@@
$f(h,\rho)=-\log(h+H(\widehat{\mathcal{G}}(\rho))-H(\widehat{\mathcal{Z}}(\rho)))-\operatorname{logdet}(\rho)$ (12)
@@
In order to implement the cone (7) in the solver Hypatia, we need to provide the gradient, Hessian, and third-order derivative of the barrier function (12), and choose an initial point [coey2023].
@@ 3.1 Gradient, Hessian, and third-order derivative
$\nabla_{h,h}^{2}f(h,\rho)$ $=\frac{1}{u^{2}}$ (22)
$\nabla_{h,\rho}^{2}f(h,\rho)$ $=\frac{1}{u^{2}}\nabla_{\rho}\,u$ (23)
@@
$\nabla_{h,h,h}^{3}f(h,\rho)$ $=-\frac{2}{u^{3}}$ (26)
$\nabla_{h,h,\rho}^{3}f(h,\rho)$ $=-\frac{2}{u^{3}}\nabla_{\rho}u$ (27)
$\nabla_{h,\rho,\rho}^{3}f(h,\rho)$ $=\xi\mapsto-\frac{2}{u^{3}}\operatorname{tr}\big{[}(\nabla_{\rho}\,u)\xi\big{]}\nabla_{\rho}u+\frac{1}{u^{2}}\nabla_{\rho,\rho}^{2}\,u[\xi]$ (28)
@@
In particular
@@
$\nabla_{\rho,\rho,\rho}^{3}u[\xi,\xi]=-{\widehat{\mathcal{G}}}^{\dagger}(U_{\mathcal{G}}M_{\mathcal{G}}(\xi)U_{\mathcal{G}}^{\dagger})+\widehat{\mathcal{Z}}^{\dagger}(U_{\mathcal{Z}}M_{\mathcal{Z}}(\xi)U_{\mathcal{Z}}^{\dagger})$ (31)
@@ 4.1 BB84
In order to illustrate the technique let us start with a toy example, the computation of the key rate for the entanglement based version of the BB84 protocol, using as estimated parameters the QBERs in the $X$ and $Z$ bases $q_{x}$ and $q_{z}$. If we use only the $Z$ basis to generate key, the key map $\mathcal{G}$ can be taken to be identity, and the decoherence map $\mathcal{Z}$ is $\rho\mapsto\sum_{i=0}^{1}(\mathopen{}\mathclose{{}\left|i\middle\rangle\middle\langle i}\right|\otimes\mathds{1})\rho(\mathopen{}\mathclose{{}\left|i\middle\rangle\middle\langle i}\right|\otimes\mathds{1})$. The analytical expression for $H(A|E)$ is $1-h(q_{x})$, where $h$ is the binary entropy.
@@
$\begin{gathered}\min_{h,\rho}\ h\\ \text{s.t.}\quad\operatorname{tr}(\rho)=1,\quad\operatorname{tr}(Q_{x}\rho)=q_{x},\quad\operatorname{tr}(Q_{z}\rho)=q_{z}\\ (h,\rho)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}},\end{gathered}$ (38)
@@ 4.5 Dealing with experimental data
$\begin{gathered}\min_{h,\sigma,\mathbf{p}}\ h\\ \text{s.t.}\quad\operatorname{tr}(\sigma)=1,\quad\operatorname{tr}(\mathbf{F}\sigma)=\mathbf{p},\quad\big{\|}\Sigma^{-\frac{1}{2}}(\mathbf{p}-\mathbf{f})\big{\|}_{2}\leq\chi\\ (h,\sigma)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}\end{gathered}$ (39)
@@ 5.1 Performance
The protocols we benchmarked are the DMCV, MUB and overlap protocols described in subsections 4.2, 4.3 and 4.4. For the DMCV protocol we used the same parameters as in [hu2022]: amplitude of the coherent states $\alpha=0.35$, noise $\xi=0.05$, distance $L=60$, and transmittance $\eta=10^{-0.02L}$. For the MUB and overlap protocols we used the probabilities from the isotropic state $\rho_{\text{iso}}(v)=v\mathopen{}\mathclose{{}\left|\phi^{+}\middle\rangle\middle\langle\phi^{+}}\right|+(1-v)\mathds{1}/d^{2}$ with visibility $v=0.95$. For the MUB protocol we computed the probabilities using all bases, included the complex ones, whereas for the overlap protocol we restricted ourselves to the real ones. We did this in order to illustrate an additional advantage of our technique: since it uses standard conic optimization, we can avail ourselves of standard symmetrization techniques to show that we can optimize over real variables only, as done in [araujo22]. This provides an additional boost in performance.
The calculations were done on an AMD Ryzen Threadripper Pro 5955WX with 4 GHz and 16 cores, on a machine with 512 GiB RAM running Ubuntu Linux 22.04. Our code was run with Julia 1.10 and the modeller JuMP [lubin2023], and the two other techniques with MATLAB 2023b. For the Gauss-Radau technique we used the solver MOSEK 10.1 [mosek] and the modeller YALMIP [yalmip]. In all cases we have only reported the time taken by the solver, discounting the time taken to set up the problem, and in the Gauss-Newton case the time taken to do facial reduction numerically.
@@ 8 Related work
While conducting this research, we found that He et al. [he2024] had independently applied the Skajaa and Ye algorithm to the problem of QKD using a specialized cone, and proved the self-concordance of the barrier function (12).
@@ 9 Acknowledgements
The research of A.G.L., P.V.P., M.C.C. and M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León. We thank Chris Coey and Lea Kapelevich for help with Hypatia’s code and useful discussions.
@@ Appendix A Vectorization
$\operatorname{vec}(ABC)=C^{T}\otimes A\operatorname{vec}(B),$ (41)
@@
$\langle X,Y\rangle=\langle\operatorname{vec}(X),\operatorname{vec}(Y)\rangle=\langle\operatorname{svec}(X),\operatorname{svec}(Y)\rangle\quad\forall X,Y$ (42)

### v3

@@ Abstract
Computing key rates in quantum key distribution (QKD) numerically is essential to unlock more powerful protocols, that use more sophisticated measurement bases or quantum systems of higher dimension. It is a difficult optimization problem, that depends on minimizing a convex non-linear function: the (quantum) relative entropy. Standard conic optimization techniques have for a long time been unable to handle the relative entropy cone, as it is a non-symmetric cone, and the standard algorithms can only handle symmetric ones. Recently, however, a practical algorithm has been discovered for optimizing over non-symmetric cones, including the relative entropy. Here we adapt this algorithm to the problem of computation of key rates, obtaining an efficient technique for lower bounding them. In comparison to previous techniques it has the advantages of flexibility, ease of use, and above all performance.
@@ 1 Introduction
We use the framework from [winick2018] to formulate the key rate problem as a convex optimization problem, the minimization of a (quantum) relative entropy over a convex domain. Using the relative entropy cone, which is a non-symmetric cone, it becomes a conic optimization problem. Although for a long time algorithms for optimizing over non-symmetric cones have been known [nesterov1999, tuncel2001, nesterov2012], they were hardly practical, as they required for instance a tractable barrier function for the dual cone. This changed with the discovery of Skajaa and Ye’s algorithm [skajaa2015, papp2017], and with it solving the key rate problem becomes in principle possible. There is, however, a technical difficulty: this formulation of the key rate problem is often not strictly feasible, and such problems cannot be reliably solved with conic optimization methods. The standard technique to deal with this, called facial reduction [drusvyatskiy2017], does apply, but the resulting reduced problem is no longer a minimization of the relative entropy. For this reason we introduce a new convex cone to deal directly with the reduced problem.
@@ 2 QKD rate as a conic problem
$\begin{gathered}\min_{\rho}\ D(\mathcal{G}(\rho)\|\mathcal{Z}(\mathcal{G}(\rho)))\\ \text{s.t.}\quad\rho\succeq 0,\quad\operatorname{tr}(\rho)=1,\\ \operatorname{tr}(E_{k}\rho)=p_{k}\ \forall k\end{gathered}$ (3)
@@
$\mathcal{K}_{\text{RE}}=\text{cl}\,\{(h,\rho,\sigma)\in\mathbb{R}\times\mathbb{H}^{n}\times\mathbb{H}^{n};\,\rho\succ 0,\ \sigma\succ 0,\ h\geq D(\rho\|\sigma)\},$ (4)
where $\mathbb{H}^{n}$ denotes the space of $n\times n$ complex Hermitian matrices. It becomes then
$\begin{gathered}\min_{h,\rho}\ h\\ \text{s.t.}\quad\rho\succeq 0,\quad\operatorname{tr}(\rho)=1,\\ \operatorname{tr}(E_{k}\rho)=p_{k}\ \forall k,\\ \Big{(}h,\mathcal{G}(\rho),\mathcal{Z}(\mathcal{G}(\rho))\Big{)}\in\mathcal{K}_{\text{RE}}\end{gathered}$ (5)
@@
$\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}=\{(h,\sigma)\in\mathbb{R}\times\mathbb{H}^{n};\,\sigma\succeq 0,\ h\geq-H(\widehat{\mathcal{G}}(\sigma))+H(\widehat{\mathcal{Z}}(\sigma))\}.$ (7)
@@
$\begin{gathered}\min_{h,\sigma}\ h\\ \text{s.t.}\quad\operatorname{tr}(\sigma)=1,\quad\operatorname{tr}(F_{k}\sigma)=p_{k}\ \forall k,\\ (h,\sigma)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}\end{gathered}$ (8)
@@ 3 Implementation of the QKD cone
In order to successfully optimize over a given convex cone $\mathcal{K}$, both using Skajaa and Ye’s algorithm or traditional methods for symmetric cones, it’s essential to provide a logarithmically homogeneous self-concordant barrier (LHSCB) function for it. It is defined as a function $f:\operatorname{int}(\mathcal{K})\to\mathbb{R}$ such that $f(x_{i})\to\infty$ along every sequence converging to the boundary of $\mathcal{K}$, is three times continuously differentiable, strictly convex, and
$f(\tau x)=f(x)-\nu\log\tau\quad\forall x\in\operatorname{int}(\mathcal{K}),\forall\tau>0,$ (9)
$|\nabla^{3}f(x)[\xi,\xi,\xi]|\leq 2(\nabla^{2}f(x)[\xi,\xi])^{3/2}\quad\forall x\in\operatorname{int}(\mathcal{K}),\forall\xi,$ (10)
where $\nabla^{2}f(x)[\xi,\xi]$ and $\nabla^{3}f(x)[\xi,\xi,\xi]$ denote the Hessian and third-order derivative at point $x$ in direction $\xi$. For an introduction to LHSCB functions and the matrix derivatives see [nesterov2018].
@@
$f(h,\rho,\sigma)=-\log(h-D(\rho\|\sigma))-\operatorname{logdet}(\rho)-\operatorname{logdet}(\sigma).$ (11)
@@
$f(h,\rho)=-\log(h+H(\widehat{\mathcal{G}}(\rho))-H(\widehat{\mathcal{Z}}(\rho)))-\operatorname{logdet}(\rho).$ (12)
@@
In order to implement the cone (7) in the solver Hypatia, we need to provide the gradient, Hessian, and third-order derivative of the barrier function (12), and choose an initial point [coey2023]. We remark that third-order derivatives are usually not required by primal-dual methods, and in particular not by the original Skajaa-Ye algorithm. Incorporating them is an innovation of Hypatia’s algorithm.
@@ 3.1 Gradient, Hessian, and third-order derivative
$\nabla_{h,h}^{2}f(h,\rho)$ $=\frac{1}{u^{2}},$ (22)
$\nabla_{h,\rho}^{2}f(h,\rho)$ $=\frac{1}{u^{2}}\nabla_{\rho}\,u,$ (23)
@@
$\nabla_{h,h,h}^{3}f(h,\rho)$ $=-\frac{2}{u^{3}},$ (26)
$\nabla_{h,h,\rho}^{3}f(h,\rho)$ $=-\frac{2}{u^{3}}\nabla_{\rho}u,$ (27)
$\nabla_{h,\rho,\rho}^{3}f(h,\rho)$ $=\xi\mapsto-\frac{2}{u^{3}}\operatorname{tr}\big{[}(\nabla_{\rho}\,u)\xi\big{]}\nabla_{\rho}u+\frac{1}{u^{2}}\nabla_{\rho,\rho}^{2}\,u[\xi],$ (28)
@@
In particular, applying $\nabla_{\rho,\rho,\rho}^{3}f(h,\rho)$ to the point $[\xi,\xi]$ gives
@@
$\nabla_{\rho,\rho,\rho}^{3}u[\xi,\xi]=-{\widehat{\mathcal{G}}}^{\dagger}(U_{\mathcal{G}}M_{\mathcal{G}}(\xi)U_{\mathcal{G}}^{\dagger})+\widehat{\mathcal{Z}}^{\dagger}(U_{\mathcal{Z}}M_{\mathcal{Z}}(\xi)U_{\mathcal{Z}}^{\dagger}),$ (31)
@@ 4.1 BB84
In order to illustrate the technique let us start with a toy example, the computation of the key rate for the entanglement based version of the BB84 protocol, using as estimated parameters the qubit error rates (QBERs) in the $X$ and $Z$ bases $q_{x}$ and $q_{z}$. If we use only the $Z$ basis to generate key, the key map $\mathcal{G}$ can be taken to be identity, and the decoherence map $\mathcal{Z}$ is $\rho\mapsto\sum_{i=0}^{1}(\mathopen{}\mathclose{{}\left|i\middle\rangle\middle\langle i}\right|\otimes\mathds{1})\rho(\mathopen{}\mathclose{{}\left|i\middle\rangle\middle\langle i}\right|\otimes\mathds{1})$. The analytical expression for $H(A|E)$ is $1-h(q_{x})$, where $h$ is the binary entropy.
@@
$\begin{gathered}\min_{h,\rho}\ h\\ \text{s.t.}\quad\operatorname{tr}(\rho)=1,\quad\operatorname{tr}(Q_{x}\rho)=q_{x},\quad\operatorname{tr}(Q_{z}\rho)=q_{z},\\ (h,\rho)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}},\end{gathered}$ (38)
@@ 4.5 Dealing with experimental data
$\begin{gathered}\min_{h,\sigma,\mathbf{p}}\ h\\ \text{s.t.}\quad\operatorname{tr}(\sigma)=1,\quad\operatorname{tr}(\mathbf{F}\sigma)=\mathbf{p},\quad\big{\|}\Sigma^{-\frac{1}{2}}(\mathbf{p}-\mathbf{f})\big{\|}_{2}\leq\chi,\\ (h,\sigma)\in\mathcal{K}_{\text{QKD}}^{\widehat{\mathcal{G}},\widehat{\mathcal{Z}}}\end{gathered}$ (39)
@@ 5.1 Performance
The protocols we benchmarked are the DMCV, MUB and overlap protocols described in subsections 4.2, 4.3 and 4.4. We chose them because they are currently of high experimental and theoretical interest [bulla2023, lib24, pascualgarcia2024, primaatmaja2024], and because they have a free parameter controlling the dimension that allows us to see how the solution time scales.
For the DMCV protocol we used the same parameters as in [hu2022]: amplitude of the coherent states $\alpha=0.35$, noise $\xi=0.05$, distance $L=60$, and transmittance $\eta=10^{-0.02L}$. For the MUB and overlap protocols we used the probabilities from the isotropic state $\rho_{\text{iso}}(v)=v\mathopen{}\mathclose{{}\left|\phi^{+}\middle\rangle\middle\langle\phi^{+}}\right|+(1-v)\mathds{1}/d^{2}$ with visibility $v=0.95$. For the MUB protocol we computed the probabilities using all bases, included the complex ones, whereas for the overlap protocol we restricted ourselves to the real ones. We did this in order to illustrate an additional advantage of our technique: since it uses standard conic optimization, we can avail ourselves of standard symmetrization techniques to show that we can optimize over real variables only, as done in [araujo22]. This provides an additional boost in performance.
The calculations were done on an AMD Ryzen Threadripper Pro 5955WX with 4 GHz and 16 cores, on a machine with 512 GiB RAM running Ubuntu Linux 22.04. Our code was run with Julia 1.11.1 and the modeller JuMP [lubin2023], and the two other techniques with MATLAB 2023b. For the Gauss-Radau technique we used the solver MOSEK 10.1 [mosek] and the modeller YALMIP [yalmip]. In all cases we have only reported the time taken by the solver, discounting the time taken to set up the problem, and in the Gauss-Newton case the time taken to do facial reduction numerically.
@@ 8 Related work
While conducting this research, we found that He et al. [he2024] had independently applied the Skajaa and Ye algorithm to the problem of QKD using a specialized cone, and proved the self-concordance of the barrier function (12). In version 0.2 of our code we implemented the technique for inverting the Hessian described in their Appendix B.2.
@@ 9 Acknowledgements
The research of A.G.L., P.V.P., M.C.C. and M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León. M.A was also supported by the Spanish Agencia Estatal de Investigación, Grant No. RYC2023-044074-I. We thank Chris Coey and Lea Kapelevich for help with Hypatia’s code and useful discussions.
@@ Appendix A Vectorization
$\operatorname{vec}(ABC)=C^{T}\otimes A\operatorname{vec}(B).$ (41)
@@
$\langle X,Y\rangle=\langle\operatorname{vec}(X),\operatorname{vec}(Y)\rangle=\langle\operatorname{svec}(X),\operatorname{svec}(Y)\rangle\quad\forall X,Y.$ (42)